OpenStudy (kitten_is_back):

Which sets of numbers are closed under subtraction? Choose all answers that are correct. A. odd natural numbers B. rational numbers C. {0, 1} D. {0, 1, 2}

OpenStudy (anonymous):

Hm... I've never heard it called "closed" under subtraction.

OpenStudy (kitten_is_back):

mjm

OpenStudy (bahrom7893):

Basically this is saying, which set can you manipulate in a way that it maintains its elements. There's an example here: http://www.algebraden.com/closure_property_subtraction_of_integers.htm

OpenStudy (bahrom7893):

Let's take C for example: {0, 1} 1 - 0 = 1 which is in the set, but 0-1 = -1 which is not. So this set is NOT closed under subtraction.

OpenStudy (bahrom7893):

or D: {0, 1, 2} again 1-2 = -1, not in the set, so not closed under subtraction

OpenStudy (kitten_is_back):

its a b

OpenStudy (bahrom7893):

What about A? What are odd natural numbers?

OpenStudy (bahrom7893):

Well, what are odd natural numbers?

OpenStudy (kitten_is_back):

idk and night '='

OpenStudy (bahrom7893):

odd natural numbers are odd positive integers: 1, 3, 5, etc. so 1-3 = -2 which is NOT an odd positive integer which is why this set is NOT closed under subtraction